OpenStudy (anonymous):

if a and b are two psoitive odd integers and if a is greater than b then prove that either a+b/2 and a-b/2 is odd and one is even

OpenStudy (anonymous):

Should we prove that (a-b)/2 is odd and (a + b)/2 is even??

OpenStudy (anonymous):

noo we have to prove any of them is even

OpenStudy (anonymous):

See, we can take a = 2n + 3 and b = 2n + 1 . Here a>b (a + b)/2 = 2n + 1 + 2n + 3/2 = 4n + 4/2 = 2n + 2 which is even (a - b)/2 = 2n + 3 - 2n - 1/2 = 2/2 = 1 which is odd Hope you understand!!

OpenStudy (anonymous):

but how bro

OpenStudy (anonymous):

See, the general form of a odd number is 2n + 1 where n is any integer. You agree or not?

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

i took the numbers as 2n + 1 and 2n + 3 . Both are odd

OpenStudy (anonymous):

i got it brother...... th nxx where r u from

OpenStudy (anonymous):

india

OpenStudy (anonymous):

nd u??